Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Test-Stand, Page 10 of 53  Not logged in ELOG logo
ID Date Author Target/No.# Source Station Module Target Oven W.O# Heat Shield W.O# Status Subject Text Attachments
  867   Tuesday, March 30, 2021, 13:19 Carla Babcock       Second emittance meter testingEmittance meter number 2 ready for beam
shortly - installed and pumping down.
No beam found by Thursday evening,
 Carla_tune_tracking.docx 
  866   Friday, March 26, 2021, 17:04 Carla Babcock       Beam line vented for emittance meter exchangeVented beamline and removed ARIEL emittance
meter. In the process of swapping components,
beamline left vented until Monday when we
  
  865   Thursday, March 25, 2021, 08:11 Aaron Schmidt       Test Stand Tripped. (Water leak)   Mar_2021_-_FS#3_leak..docx 
  864   Monday, March 15, 2021, 16:10 Carla Babcock       TISA emittance meter testing with new FEBIAD sourceMonday, March 15, 2021, 16:09 Target and
ion source ramped to 400A/330A this afternoon
Monday, March 15, 2021, 18:53 : Tune
 Carla_tune_tracking.docx 
  863   Wednesday, March 10, 2021, 14:57 Aaron Schmidt HP-FEBIAD   42752-5In progressEmittance meter test with FEBIADWednesday, March 10, 2021, 14:54: Roughing.
Wednesday, March 10, 2021, 14:57:
BV1A open.
  
  862   Wednesday, March 10, 2021, 14:28 Aaron Schmidt      In progressRP2 swap.Swapped RP2 out from a Varian SH-110 to
a Varian PTS06001.
Wednesday, March 10, 2021, 14:20:
  
  861   Wednesday, March 10, 2021, 11:10 Carla Babcock       Venting for installation of emittance metersWednesday, March 10, 2021, 10:48 
: Venting beamline section only for installation
of emittance meters.
  
  860   Tuesday, March 02, 2021, 11:42 Aaron Schmidt       TS Tripped.    
  859   Monday, March 01, 2021, 16:13 Fernando Maldonado      DoneTemporary RGA installation Monday, March 01, 2021, 16:03: An
ExTorr RGA has been installed in the ISAC
test stand.
 EXTorr.PNG 
  858   Monday, March 01, 2021, 12:58 Fernando Maldonado LP-FEBIAD    DoneFEBIAD parameter scans v2 Monday, March 01, 2021, 12:57: Roughing
Monday, March 01, 2021, 13:30: Turbos
on. Started a bit late as I was distracted.
  
  857   Monday, February 22, 2021, 18:25 Fernando Maldonado HP-FEBIAD    DoneFEBIAD parameter scansMonday, February 22, 2021, 17:48 Roughing
Monday, February 22, 2021, 18:17:
Turbos on.
 ESC_-_Feb_23_FEBAID_parameter_initial_checks.docxER-580-290-30kV-m28_2021-2-23.txt 
  856   Friday, February 19, 2021, 13:44 Fernando Maldonado      FailedFEBIAD parameter scansFriday, February 19, 2021, 13:28: Roughing
Friday, February 19, 2021, 13:43:
Power on
 GND-COIL_spark.jpgGND-HS_spark.jpg 
  855   Thursday, February 18, 2021, 16:48 Fernando Maldonado      FailedFEBIAD parameter scansThursday, February 18, 2021, 15:19: Gas
lines roughed and refilled. Gauge reads -50
kPa. Absolute pressure is 50 kPa.
  
  854   Thursday, February 18, 2021, 15:15 Fernando Maldonado      DoneHLA testsThursday, February 18, 2021, 15:01: With
the help of Spencer Kiy, we tested the proper
communication from python to epics.
  
  853   Thursday, February 18, 2021, 11:03 Aaron Schmidt      Problems/SolutionsDI water pump configurationhttps://elog.triumf.ca/TIS/Test-Stand/847
There was a previous DI CU-LCW 
pump trip as per the above elog. The CU-LCW
  
  852   Thursday, February 18, 2021, 10:47 Aaron Schmidt       Optics loss / CANBUS failure.Thursday, February 18, 2021, 10:45: TPS:Q3
& the TSB CBC CANBUS controllers had
failed yesterday, as well as a cable was
  
  851   Tuesday, February 16, 2021, 14:52 Fernando Maldonado      Problems/SolutionsLeak checking gas lines.Vacuum side
Tuesday, February 16, 2021, 13:42:
The gas lines inside have been check. The
  
  850   Friday, February 12, 2021, 16:38 Fernando Maldonado      FailedFEBIAD parameter scan - efficiency measurements - energy spread.Friday, February 12, 2021, 16:37: FEBIAD
installed with mechanical calibrated leak
Friday, February 12, 2021, 16:38:
  
  849   Friday, February 12, 2021, 12:16 Fernando Maldonado      Problems/SolutionsGrounding collimator cable from outside Friday, February 12, 2021, 12:13:
Since the collimator is not used, Carla and
I installed a BNC short outside the cage.
 20210212_121227.jpg 
  848   Wednesday, February 10, 2021, 09:49 Aaron Schmidt LP-SIS  44691-844378-2 / 44378-3FailedEnergy spread calibration with SISNote* target has a ring on the TGHT - Right
leg. EE & PE placed. Target to HS OL
- OK.
 spark_inside_chamber_roof.pngspark_inside_chamber_front_view.png 
ELOG V2.9.2-2455